# Plasmons in two-dimensional electron systems with infinite and semi-infinite metal gratings

### v1

@@ I Introduction
The present work is focused on plasma oscillations in a 2DES with an infinite or semi-infinite periodic array (grating) of metal strip gates close to the 2DES, as shown in Fig. 1. Plasmons in such systems have been studied in sufficient detail in the literature [33, 35, 36, 10, 11, 37, 19, 38, 39, 40, 41, 42, 43]. However, to the best of our knowledge, these studies considered only the plasma waves propagating across the strip-shaped gates (in $x$-direction in Fig. 1). Therefore, this paper also examines the plasmons propagating along the strip gates, i.e. with a nonzero wave vector $q_{y}$, as designated in Fig. 1.
@@ III Analysis of plasmons in the 2DES with infinite and semi-infinite gratings
Fig. 2 depicts the first and second frequency bands obtained according to (8), where the dark-green lines indicate the boundaries of the bands defined by $f(q_{y},\omega)=\pm 1$. In the figure, we can see that in the short-wavelength limit of $|q_{y}|b\gg 1$, when the wave vector $q_{y}$ or the distance between the gates $b$ increases, the width of plasmonic bands decreases and becomes infinitely small. At the same time, the boundaries of the band of number $K$ tend asymptotically to the relation for the gated plasmon $\omega=V_{p}\sqrt{\pi^{2}K^{2}/W^{2}+q_{y}^{2}}$.
@@ IV Discussion and conclusion
It should be mentioned as well that in the framework of the optical approach, the interaction of 2D electrons with their mirror-images induced in the respective gates is considered to be dominant, while the direct interaction between 2D electrons under the gates and the interaction of electrons in the gates are regarded negligible. For this reason, the distance between the gates, $b$, is assumed to be much larger than the distance $d$ separating the 2DES and the grating.
Finally, in the derivation of dispersion equations (7), (8), and (9), we neglect the plasmon charge density excited in the ungated regions of the 2DES, taking the right-hand side of Eq. (5) to be equal to zero. It is justified by the absence of ungated plasmons in the system since their frequency far exceeds that of the plasmons, that are considered in this paper, at the same wave vector $q_{y}$. Namely, the frequency of ungated plasmons with the wave vector $q_{y}$ (at $q_{y}d\ll 1$) is given as [1]:
@@
By relative comparison to the characteristic frequency of the Tamm plasmon, for example, using Eq. (19), we find that $\omega_{Tamm}(q_{y})/\omega_{p}(q_{y})=\sqrt{2d(\varepsilon+1)/(\varepsilon W_{0})}$, which is much less than unity, as $d\ll W_{0}$ in our study. In general, if $\omega_{p}(q_{y})$ is greater than the frequency of the Tamm plasmons (for instance, in the case of large $N$), it is likely that they decay into the continuum of ungated plasmons existing in the 2DES at $x<-W_{0}$ (Fig. 1(b)). Indeed, such decay has been considered in Ref. [48] for plasmons in a homogeneous 2DES with a disk-shaped gate.
@@ Appendix B Derivation of the equations for infinite and semi-infinite grating gates
which equals to zero according to Eq. (29). At the same time, the real part can be expressed as:

### v2

@@ I Introduction
The present work is focused on plasma oscillations in a 2DES with an infinite or semi-infinite periodic array (grating) of metal strip gates close to the 2DES, as shown in Fig. 1. Plasmons in such systems have been studied in sufficient detail in the literature [33, 35, 36, 10, 11, 37, 19, 38, 39, 40, 41, 42, 43, 44]. However, to the best of our knowledge, these studies considered only the plasma waves propagating across the strip-shaped gates (in $x$-direction in Fig. 1). Therefore, this paper also examines the plasmons propagating along the strip gates, i.e. with a nonzero wave vector $q_{y}$, as designated in Fig. 1.
@@ III Analysis of plasmons in the 2DES with infinite and semi-infinite gratings
Fig. 2 depicts the first and second frequency bands obtained according to (8), where the dark-green lines indicate the boundaries of the bands defined by $f(q_{y},\omega)=\pm 1$. In the figure, we can see that in the short-wavelength limit of $|q_{y}|b\gg 1$, when the wave vector $q_{y}$ or the distance between the gates $b$ increases, the width of plasmonic bands decreases and becomes infinitely small. At the same time, the boundaries of the band of number $K$ tend asymptotically to the relation for the gated plasmon $\omega=V_{p}\sqrt{\pi^{2}K^{2}/W^{2}+q_{y}^{2}}$. Let us emphasize that the formation of the plasmonic band structure in our study is due to the interaction of currents (and electric fields) flowing in ungated parts of 2DES. These fields and currents are generated by oscillated charge density in gates as well as gated areas of 2DES and can extend into ungated regions over distances up to $q_{y}^{-1}$, while the charge density in ungated regions is neglected, see Eq. (5).
@@ IV Discussion and conclusion
It should be mentioned as well that in the framework of the optical approach, the interaction of 2D electrons with their mirror-images induced in the respective gates is considered to be dominant, while the interaction between electrons in different gated region of 2DES and in different gates is regarded negligible. For this reason, the distance between the gates, $b$, is assumed to be much larger than the distance $d$ separating the 2DES and the grating.
Also, in the derivation of dispersion equations (7), (8), and (9), we neglect the plasmon charge density excited in the ungated regions of the 2DES, taking the right-hand side of Eq. (5) to be equal to zero. The charge density in ungated regions of 2DES can appear via two main reasons. The first one is the excitation of ungated plasmons, however, such plasmons are absent in the system since their frequency far exceeds that of the plasmons, that are considered in this paper, at the same wave vector $q_{y}$. Namely, the frequency of ungated plasmons with the wave vector $q_{y}$ is given as [1]:
@@
By relative comparison to the characteristic frequency of the Tamm plasmon, for example, using Eq. (19), we find that $\omega_{Tamm}(q_{y})/\omega_{p}(q_{y})=\sqrt{2d(\varepsilon+1)/(\varepsilon W_{0})}$, which is much less than unity, as $d\ll W_{0}$ in our study. In general, if $\omega_{p}(q_{y})$ is greater than the frequency of the Tamm plasmons (for instance, in the case of large $N$), it is likely that they decay into the continuum of ungated plasmons existing in the 2DES at $x<-W_{0}$ (Fig. 1(b)). Indeed, such decay has been considered in Ref. [49] for plasmons in a homogeneous 2DES with a disk-shaped gate. The second main reason of the appearance of nonzero charge density in ungated regions of 2DES is the ‘induction’ of this charge by the charge in the adjacent gate. The value of this induced in ungated region of 2DES charge can be qualitatively characterized from the estimation of the capacitance between the 2DES and the gate. Indeed, as shown, for instance, in Appendix of Ref. [42] the capacitance has two contributions: the first one stems from the interaction of electrons in the gate with 2D electrons under the gate, while the second one corresponds to the interaction with 2D electrons in ungated area. One can see from Eq. (A5) of Ref. [42] that the first term is larger than the second one by the parameter $W/d$. Qualitatively, this can be understood by the fact the charge in the ungated part of 2DES is induced at distances not exceeding several $d$, see also Figs. 4(a) and (b) from Ref. [45]. Thus, in the case of our system the charge in the ungated regions of the 2DES can be neglected.
@@
Finally, let us compare briefly the obtained results with previous studies of Tamm plasmons in 2DES with grating gate performed in Ref. [60], in which the dispersion relation for plasmons propagating perpendicular to gates was derived, see Eq. (23) from Ref. [60], via the introduction of electrical impedances for different regions of 2DES. Using local capacity approximation ($qd\ll 1$) and setting $q_{2}=0$, Eq. (23) from Ref. [60] becomes the same as Eq. (7), in which $q_{y}$ should be taken as zero. As for Tamm modes, the spectrum found in this paper at $q_{y}=0$ qualitatively matches with the spectrum obtained in Ref. [60] for the case of equal electron concentrations in gated and ungated parts of 2DES.
@@ Appendix B Derivation of the equations for infinite and semi-infinite grating gates
which equals zero according to Eq. (29). At the same time, the real part can be expressed as: